if a father is 43 years old and his son is 7 years old, in how many years will the father be three times as old as the son?
Answer by checkley79(3341)

(45+X)/(7+X)=3/1
3(7+X)=45+X
21+3X=45+X
3X-X=45-21
2X=24
X=24/2
X=12 YEARS FROM NOW IS THE ANSWER.
PROOF:
(45+12)/(7+12)=3/1
57/19=3/1
3*19=1*57
57=57